OpenStudy (anonymous):

how many solutions are there? 9x+3y=−6 and y=−x+5

OpenStudy (anonymous):

so it's like this 9x+3y=-6 y=-x+5 9x+3*(-x+5)=-6 9x+3*(-x)+3*5=-6 9x-3x+15=-6 6x+15=-6 6x=-6-15 6x=-21 x=-21/6 x=-3,5 there is only ONE solution